whitemonsterbunny17 (whitemonsterbunny17):

Solve for x. 4(x-6)-6x=-18

whitemonsterbunny17 (whitemonsterbunny17):

I don't want someone to solve this for me; I just want someone to show me how to do this type of problem.

whitemonsterbunny17 (whitemonsterbunny17):

@ganeshie8

whitemonsterbunny17 (whitemonsterbunny17):

@horsegirl27

OpenStudy (horsegirl27):

You distribute and combine like terms

OpenStudy (horsegirl27):

Then simplify

whitemonsterbunny17 (whitemonsterbunny17):

So... x-6+6x?

whitemonsterbunny17 (whitemonsterbunny17):

I have no idea... I haven't done algebra in about 4-5 years...

whitemonsterbunny17 (whitemonsterbunny17):

OR wait.. I distribute the 4, right?

whitemonsterbunny17 (whitemonsterbunny17):

So 4(x-6*4)-6x*4=-18*4 ?

OpenStudy (horsegirl27):

No 4 into x and -6

OpenStudy (horsegirl27):

so 4x-24-6x=-18 Now combine all the x's

whitemonsterbunny17 (whitemonsterbunny17):

4x+(-6x)?

OpenStudy (horsegirl27):

Yup that's -2x=-18 You know what to do next?

whitemonsterbunny17 (whitemonsterbunny17):

Take -18+(-2x)?

whitemonsterbunny17 (whitemonsterbunny17):

or divide?

whitemonsterbunny17 (whitemonsterbunny17):

Yes, divide.

whitemonsterbunny17 (whitemonsterbunny17):

Which equals -9, right?

OpenStudy (horsegirl27):

Good job!

whitemonsterbunny17 (whitemonsterbunny17):

Wait, I was right? :o YAY! :D

OpenStudy (horsegirl27):

Wait hold that thought, it's positive 9

whitemonsterbunny17 (whitemonsterbunny17):

Oh. :p I was close. xD

whitemonsterbunny17 (whitemonsterbunny17):

Could you help me with another? It's an inequality one. :p

whitemonsterbunny17 (whitemonsterbunny17):

Also, thanks for putting up with me, I know I'm difficult. xD You're awesome though! ^.^

OpenStudy (horsegirl27):

Haha it's okay, yo're welcome